### v1

Abstract
Understanding the dynamics of magnons in quantum magnetism is a captivating area of condensed matter physics. Magnons exhibit unique collective behaviours, including the formation of multimagnon bound states (MBSs). Predicted by Bethe over 90 years ago, MBSs in one-dimensional (1D) quantum magnets remain a significant subject of study due to their fundamental and technological implications. While recent advancements in quantum simulation with ultracold atoms have enabled the observation of MBSs in optical lattices, observing these states in naturally occurring materials presents an even more exciting prospect.
This study proposes a novel mechanism for the stabilisation of MBSs in quasi-1D (Q1D) edge-shared cuprates, materials known for their complex magnetic interactions. Contrary to conventional approaches that require high external magnetic fields, our mechanism leverages small antiferromagnetic (AFM) interchain couplings, which act analogously to an internal magnetic field. These couplings induce collinear AFM order within individual chains, facilitating the formation of stable MBSs. This self-organised magnon condensation driven by interchain interactions is experimentally more accessible and opens new avenues for observations.
Our findings offer a fresh perspective on magnetic ordered states driven by interchain-coupling-induced internal magnetic fields. We provide a comprehensive theoretical framework, present numerical results compared with experimental data from specific Q1D cuprates like Li2CuO2, Ca2Y2Cu5O10, LiCuSbO4, and PbCuSO4(OH)2, and discuss the broader implications of our discoveries. This work paves the way for future research exploring the potential of interchain coupling-driven magnon condensation, significantly advancing the fields of quantum computing, spintronics, and magnonics.
keywords:
magnon bound state, internal field, quasi-one-dimensional cuprates, density-matrix renormalization group
1 Introduction
Understanding the dynamics of magnons in quantum magnetism is a captivating area of condensed matter physics. Magnons, the quasiparticles associated with spin waves in magnetically ordered systems, exhibit unique collective behaviours, including the formation of multimagnon bound states (MBSs). The existence of MBSs in one-dimensional (1D) quantum magnets was predicted over 90 years ago by Bethe [1]: Due to ferromagnetic (FM) interactions, two spin excitations can remain bound together, forming what is known as a two-magnon bound state [1, 2, 3]. Investigating the formation and behaviour of MBSs addresses fundamental questions in physics and holds significant potential for advancing quantum computing [4], spintronics [5], and other next-generation technologies like magnonics [6].
Recent advancements in quantum simulation with ultracold atoms have enabled the observation of magnon excitations and their two-magnon bound states in quantum Heisenberg chains comprising bosonic atoms in optical lattice experiments [7, 8]. Nonetheless, observing MBSs in naturally occurring materials would be even more exciting. Pursuing this possibility, theoretical studies have extensively investigated the emergence of MBSs in frustrated ferromagnetic-antiferromagnetic (FM-AFM) $J_{1}$-$J_{2}$ chain systems, which are typical magnetic models for quasi-1D (Q1D) edge-shared cuprates [9]. These studies have demonstrated that MBSs can be induced by a magnetic field, and interestingly, near saturation, the types of MBSs vary with the ratio of $J_{1}$ to $J_{2}$, transitioning between two-magnon (nematic), three-magnon (triatic), four-magnon (quartic), and so on [10, 11, 12, 13, 14, 15, 16, 17]. Subsequent experimental efforts have aimed to identify MBSs, focusing on highly 1D materials such as LiVCuO4, where signatures of nematic states have been observed [18, 19, 20, 21, 22, 23]. Nevertheless, the stabilisation of MBSs generally requires high external magnetic fields, making experimental observation challenging due to their fragility against thermal fluctuations and external perturbations [24, 25].
In this study, we propose a novel mechanism for the stabilisation of MBSs in Q1D edge-shared cuprates, a class of materials known for their complex magnetic interactions. Unlike conventional approaches, our mechanism does not rely on external magnetic fields for the emergence of MBSs. Instead, small AFM interchain couplings act analogously to an internal magnetic field. These interchain couplings induce a collinear AFM order (CAFO) within individual chains, leading to the formation of stable MBSs. This self-organised condensation of magnons driven by interchain interactions is much more accessible for experimental verification, opening in this way new avenues for observations. Although this perspective has been theoretically suggested, it has been somewhat overlooked when considering real materials. Thus, our findings provide a fresh perspective on revisiting various magnetically ordered states as driven by interchain-coupling-induced effective magnetic fields – we call it internal fields.
This paper is structured as follows: we first outline the theoretical framework for MBS and the role of AFM interchain couplings. Next, we present our numerical results and their comparison with experimental data from specific Q1D cuprates. Finally, we discuss the broader implications of our findings and propose future research directions to further explore the potential of interchain coupling-driven magnon condensation.
2 Magnon bound states by external field
Figure 1: (a) Schematic illustration of the MBS driven by interchain coupling, corresponding to a triatic state with three magnons bound together. The yellow lines represent transverse spin fluctuations. Crystal structures of (b) Li2CuO2, showing two CuO2 chains per unit cell along the b axis, and (c) Ca2Y2Cu5555O10, showing two CuO2 chains per unit cell along the a axis. Red and blue arrows in (b) and (c) indicate an example of symmetry-broken magnetization directions in the CAFO state. Structure of interchain couplings for compounds such as (d) Li2CuO2 and Ca2Y2Cu5555O10, and (e) LiCuSbO4 and PbCuSO4(OH)2. (f) Schematic of the chain positions used in the numerical 2-chain, 4-chain, and 8-chain calculations, where the dashed lines denote the interchain coupling (also, see Ref. [26]).
Prior to discussing our novel emergence mechanism of MBSs, it is necessary to outline the concept of MBSs and to explain how to theoretically calculate the magnon binding energy, which is one of the most direct physical quantities for detecting MBS. To illustrate the conventional MBS that emerges when spin-rotation symmetry is broken by an external field, we utilise the single FM-AFM $J_{1}$-$J_{2}$ chain. This analysis elucidates the key properties of MBSs, including the magnon binding energy as a function of the ratio between $J_{1}$ and $J_{2}$. These descriptions provide essential information for understanding the internal-field-induced MBSs discussed later.
2.1 Magnon binding
Magnons are collective excitations that arise from quantised spin wave perturbations in magnetically ordered systems, including ferromagnets, antiferromagnets and helimagnets. These entities adhere to Bose-Einstein statistics, classifying them as bosons. The presence of a magnon signifies a flip in the spin orientation of an electron relative to its ordered (ground) state. It is notable that in environments with strong magnon-magnon interactions and enhanced magnetic correlations, which are typically observed in low-dimensional systems prone to magnetic frustration, magnons exhibit unique behaviours, including the ability to pair and form MBSs. Such pairing results in a composite excitation known as a bimagnon, analogous to Cooper pairs in superconductors.
The quantum phase referred to as a spin nematic state, i.e., a two-magnon bound state, is characterised by a predominance of collective bimagnon excitations. It features a preferred orientation of the paired magnons without spatial confinement, breaking spin-rotational symmetry while retaining translational symmetry. This phenomenon is typically realised in conditions where spins are polarized, for instance, by an external magnetic field. Furthermore, under specific conditions, higher-order bound states involving more than two magnons, such as ‘triatic’ or ‘quartic’ states, may form. These states arise from the intricate interactions of three or four magnons, respectively.
An effective theoretical approach to identifying MBS is to calculate the binding energy of magnons, which is crucial for understanding their interactions and properties. This energy quantifies the cost of dissociating a bound state into individual magnons. This method is analogous to the binding strength of Cooper pairs in superconductivity [27]. When a MBS forms, it is associated with a specific binding energy. To calculate this for a $p$-MBS (where $p$ is the number of magnons, e.g. 2 for nematic, 3 for triatic, 4 for quartic state) at a given net magnetisation $M$ (assuming conserved $S^{z}$), the formula is
$E_{b}(M,p)\!=\left[E_{0}(S^{z}\!=\!M\!-\!1)\!-\!E_{0}(S^{z}\!=\!M)\right]\!-\!\frac{1}{p}\left[E_{0}(S^{z}\!=\!M\!-\!p)\!-\!E_{0}(S^{z}\!=\!M)\right],$ (1)
where the ground state energy for a total spin $S^{z}=S$ is denoted by $E_{0}(S^{z}=S)$. A positive value of the binding energy in the bulk limit indicates the formation of a bound state, which is an important indicator of the presence of MBS. However, unlike the electron binding energy in superconductivity, the number of bound magnons is not known in advance, so it is necessary to calculate this quantity for different values of $p$. If the largest value of $E_{b}(M,p)$ ($>0$) is given by $p=p_{\rm max}$, it can be concluded that the $p_{\rm max}$ MBS is more energetically favourable than both the individual magnons and other $p$ MBSs. A schematic illustration of the nematic case ($p=2$) is shown in Fig. 2a. The magnon binding energy is observed in experiments as a kind of excitation gap. The following subsection illustrates the calculation of the magnon binding energy for the FM-AFM $J_{1}$-$J_{2}$ chain.
2.2 1D frustrated chain at high fields
Figure 2: (a) Schematic illustrations of (top) the fully polarized state, (middle) the state with two flipped spins far apart – individual magnons –, and (bottom) the state with two adjacent flipped spins – bounded magnons –. (b) Propagation number $\theta$ (black circles) of the $J_{1}$-$J_{2}$ chain at zero magnetic field as a function of $J_{2}/\lvert J_{1}\rvert$, and the relationship between the number of bound magnons $p$ in the MES at the saturation field. The propagation number is fitted by $\theta/\pi=0.69(J_{2}/\lvert J_{1}\rvert)^{0.29}$ near $J_{2}/\lvert J_{1}\rvert=1/4$. (c) Magnon binding energy $E_{b}(M_{\rm s},p)$ corresponding to $p$ in the MES at the saturation field. The dashed line denotes $E_{b}(M_{\rm s},p)\propto(J_{2}/\lvert J_{1}\rvert)^{\pi/2}$. (d) Ground state phase diagram for MBS as a function of $J_{1}/J_{2}$ (and $J_{2}/\lvert J_{1}\rvert$) and $M/M_{\rm s}$, with the magnon binding energy $E_{b}(M,p)$ depicted in a colour density plot corresponding to each phase.
One of the most promising systems exhibiting MBS is the 1D Heisenberg chain with frustrated FM nearest neighbour (NN) and AFM next nearest neighbour (NNN) interactions, namely the FM-AFM $J_{1}$-$J_{2}$ chain, under an external magnetic field. The Hamiltonian for this system is expressed as follows
${\cal H}_{\alpha}=J_{1}\sum_{i}\vec{S}_{\alpha,i}\cdot\vec{S}_{\alpha,i+1}+J_{2}\sum_{i}\vec{S}_{\alpha,i}\cdot\vec{S}_{\alpha,i+2}-h\sum_{i}S^{z}_{\alpha,i},$ (2)
[…]
The ground state of the system (2) at zero field ($h=0$) has been studied extensively [28, 29, 30, 31, 32, 33, 34]. There are two distinct phases: the FM phase at $J_{2}/\lvert J_{1}\rvert<1/4$ and the incommensurate spiral phase at $J_{2}/\lvert J_{1}\rvert>1/4$. Throughout the spiral phase, the incommensurate spin-spin correlations are short-range [30, 32]; instead, a spontaneous NN-FM dimerisation order accompanied by translational symmetry breaking occurs due to the order-by-disorder mechanism [33, 34]. Consequently, the system exhibits a form of valence bond solid (VBS) state with spin singlet formations between third-neighbour sites [34]. By considering the ferromagnetically dimerised spin-1/2 pair as a spin-1 site, this VBS state can also be recognised as an Affleck-Kennedy-Lieb-Tasaki [35] or Haldane [36, 37] state with topological order protected by global $Z_{2}\times Z_{2}$ symmetry [38].
The theoretical phase diagram at finite magnetic fields has attracted considerable attention in the context of MBSs [10, 11, 12, 13, 14, 15]. When an external magnetic field is applied to the spiral phase, a vector chiral phase appears. This phase is characterised by the breaking of the discrete parity symmetry ($Z_{2}$ symmetry breaking) accompanied by a spontaneous $S^{z}$ spin current circulation [14]. The longitudinal vector chiral order parameter, $\kappa_{l}^{(n)}=\langle(\mbox{\boldmath$S$}_{l}\times\mbox{\boldmath$S$}_{l+n})^{z}\rangle$, acquires a non-zero value in this phase. As the magnetic field strength increases, an MBS appears. The number of bound magnons depends on the ratio of $J_{2}$ to $\lvert J_{1}\rvert$. By calculating the binding energy $E_{\rm b}(M,p)$, expressed in eq. 1, the number of bound magnons can be determined. Fig. 2b shows the propagation number $\theta$ at zero field and the number of bound magnons $p$ near full saturation ($M/M_{\rm s}=1$) as a function of $J_{2}/\lvert J_{1}\rvert-1/4$. The corresponding binding energy is plotted in Fig. 2c. Approaching the FM phase transition at $J_{2}/\lvert J_{1}\rvert=1/4$, the number of bound magnons increases while the binding energy decreases sharply, indicating that larger bound magnon sizes are more susceptible to external perturbations. The conjectured relationship between $\theta$ and $p$, $1/p>\theta/\pi>1/(p+1)$ [15], holds over a wide range of $J_{2}/\lvert J_{1}\rvert$ values, except near $J_{2}/\lvert J_{1}\rvert=1/4$. Furthermore, Fig. 2d shows the distribution of $E_{\rm b}(M,p)$ as a function of $J_{1}/J_{2}$ and $M/M_{\rm s}$. We observe that the binding energy peaks slightly below the saturation magnetisation $M/M_{\rm s}=1$, suggesting that sufficient breaking of the spin-rotation symmetry allows bound magnons to move parallel or antiparallel to the magnetic field, thereby gaining propagation energy.
2.3 Experimental Challenges
The experimental observation of MBS in Q1D systems presents significant difficulties. MBS can be detected by a variety of experimental methods, including the temperature dependence [39, 40, 23] and field dependence [16, 22] of the nuclear magnetic resonance (NMR) relaxation rate, field dependent wave vector measurements by neutron scattering experiments [12, 13, 14, 15, 35, 19] and field dependent magnetisation [16, 18]. Fig. 2d suggests that the MBS phase spans a wide range of magnetisation. However, the substantial quantum fluctuations in Q1D systems result in a very gradual increase in magnetisation at low fields. Thus, high magnetic fields are required to achieve sufficient magnetisation to stabilise MBS in these systems. Furthermore, while the binding energy of magnons in an isolated $J_{1}$-$J_{2}$ chain is already small, it decreases sharply with increasing AFM interchain interactions. As a result, the MBS is extremely sensitive to these AFM interchain interactions [24, 25]. Therefore, experimental setups often require conditions such as low temperatures and high magnetic fields to stabilise and observe MBS. Consequently, the practical observation of external field induced MBS remains a major challenge.
3 Magnon bound states by internal field
As previously discussed, the observation of MBS in Q1D systems has traditionally been a matter of contention, primarily due to the challenges in experimental detection. However, the novel mechanism proposed herein facilitates magnon binding via AFM interchain coupling, which can mimic the effect of an actual external magnetic field for each chain, thereby eliminating the need for an external magnetic field. This advancement presents a significant advantage for experimental observations. This section elucidates the underlying mechanism and demonstrates that MBS can indeed emerge within realistic parameter regimes.
3.1 Internal field from mean-field approximation of interchain couplings
In Q1D magnetic systems, the interplay between interchain and intrachain couplings plays a pivotal role in determining the magnetic ground state. Notably, when interchain couplings are significantly weaker than their intrachain counterparts, the interchain interactions can be effectively approximated through a mean-field approach. This approximation allows for the conceptualisation of the system as a collection of independent 1D chains influenced by an effective mean field. This concept was initially validated for isotropic, i.e., XXX, Heisenberg chains linked by perpendicular interchain couplings $J_{\perp}$, described by the Hamiltonian ${\cal H}=\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i,j+1}+J_{\perp}\sum_{i,j}\mbox{\boldmath$S$}_{i,j}\cdot\mbox{\boldmath$S$}_{i+1,j}$ with $J_{\perp}>0$. In the case of $J_{\perp}=0$, the system lacks magnetic order and is classified as a Tomonaga-Luttinger liquid at zero temperature. However, the spin-spin correlation functions decay in a power law manner, indicating the presence of quasi-long-range order. Hence, if finite $J_{\perp}$ is switched on, the system exhibits Néel long-range order [41].
Assuming the order to be oriented along the $z$-direction in spin space, the Hamiltonian transforms into an effective single-chain problem described by ${\cal H}_{\rm eff}=J\sum_{i}\mbox{\boldmath$S$}_{i}\cdot\mbox{\boldmath$S$}_{i+1}-h_{\rm stag}\sum_{i}(-1)^{i}S^{z}_{i}+{\rm const.}$, where $h_{\rm stag}$ denotes the staggered field. This field, acting as an effective internal force, encapsulates the collective effect of the interactions with adjacent chains, inherent to the structure of the system. It elucidates the formation of Néel order in 2D square and 3D cubic lattices as a dimensional extension from a 1D chain. The configuration of this effective internal field depends on the structural arrangement of magnetic interactions within the system. For example, it can exhibit incommensurate oscillations if the net magnetisation of the system is nonzero [42].
What are the potential forms of internal fields in Q1D FM-AFM $J_{1}$-$J_{2}$ chain systems? The Hamiltonian, incorporating interchain coupling, can be expressed as
${\cal H}=$ $\underbrace{J_{1}\sum_{\alpha,i}[\Delta_{1}(S^{+}_{\alpha,i}S^{-}_{\alpha,i+1}+S^{-}_{\alpha,i}S^{+}_{\alpha,i+1})+S^{z}_{\alpha,i}S^{z}_{\alpha,i+1}]+J_{2}\sum_{\alpha,i}\vec{S}_{\alpha,i}\cdot\vec{S}_{\alpha,i+2}}_{\text{intrachain couplings}}$
$+\underbrace{\sum_{\begin{subarray}{c}\alpha\beta,ij\\ \gamma=a,b\,{\rm or}\,c,d\end{subarray}}J_{\gamma}[\Delta_{\gamma}(S^{+}_{\alpha,i}S^{-}_{\beta,j}+S^{-}_{\alpha,i}S^{+}_{\beta,j})+S^{z}_{\alpha,i}S^{z}_{\beta,j}]}_{\text{interchain couplings}}$ (3)
where $J_{\gamma}$ ($\gamma={\rm a,b\,or\,c,d}$) denotes the interchain coupling, the common structures of which are illustrated in Fig. 1d,e. For example, compounds corresponding to Fig. 1d include Li2CuO2 and Ca2Y2Cu5555O10, while those corresponding to Fig. 1e include LiCuSbO4, PbCuSO4(OH)2, and LiVCuO4. In the direction perpendicular to the chain, they have a bipartite structure. If one could neglect the effects of interchain coupling, each chain would exhibit a spiral state with incommensurate short-ranged spin-spin correlations if $J_{2}/\lvert J_{1}\rvert>1/4$. The pitch angle $\phi$ of this spiral, dependent of $J_{2}/\lvert J_{1}\rvert$, ranges between 0 and $\pi/2$, and $\phi=\cos^{-1}[-J_{1}/(4J_{2})]$ in the classical limit. When the interchain couplings are AFM, it is known that the pitch angle decreases as their coupling strength increases, potentially leading to FM polarisation of each chain when $\phi$ reaches zero [43]. This indicates that the ratio $J_{2}/\lvert J_{1}\rvert$ is effectively diminished by AFM interchain coupling, reaching an FM critical point $1/4$ at a certain threshold. Conversely, it can also be said that AFM interchain coupling shifts the FM critical point towards higher $J_{2}/\lvert J_{1}\rvert$ values. The effective FM critical point can be analytically calculated, yielding for the cases represented in Fig. 1d and 1e,
$\left(\frac{J_{2}}{\lvert J_{1}\rvert}\right)_{\rm FM,c}=\frac{1}{4}+\frac{1}{8}\frac{J_{a}+9J_{b}}{\lvert J_{1}\rvert}$ (4)
and
$\left(\frac{J_{2}}{\lvert J_{1}\rvert}\right)_{\rm FM,c}=\frac{1}{4}+\frac{J_{d}}{2\lvert J_{1}\rvert},$ (5)
respectively. Note that this analysis is for a 2D system, where each chain has two adjacent chains. For a 3D system, where each chain has four adjacent chains, these equations are valid with the replacements $J_{a}\to 2J_{a}$, $J_{b}\to 2J_{b}$, and $J_{d}\to 2J_{d}$ [43].
If the magnitude of the AFM interchain coupling is sufficiently large to induce FM ordering in each chain, the spins in adjacent chains will be polarised in opposite directions. This scenario represents a phase transition from a spiral state to a CAFO, governed by the AFM interchain coupling. By conceptualising the AFM interchain coupling as an effective internal field, each chain can be viewed as an FM-AFM $J_{1}$-$J_{2}$ chain under a uniform field, with the field direction being opposite in adjacent chains. If this mapping holds true, the effective Hamiltonian for a single chain should be expressible in a form similar to Equation (2). In other words, if the internal field induced by the AFM interchain coupling can break the spin symmetry of each chain, leading to a finite magnetisation, it might be possible to realise MBSs without the need for an external magnetic field. In the following subsections, we will numerically verify the formation of CAFO as the AFM interchain coupling is increased, and the realisation of MBSs.
3.2 Magnetisation in self-organized collinear antiferromagnetic ordering
Figure 3: Magnetization per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy, evaluated at representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. The magnetization is presented as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$. Panels (a) and (b) illustrate the effects of introducing XXZ anisotropy on inchain and interchain couplings, respectively. The parameters used in the calculations of MES are marked with circles (see Fig. 4).
Here, we demonstrate the formation of CAFO driven by AFM interchain interaction in the coupled FM-AFM $J_{1}$-$J_{2}$ chains. To align the magnetisation direction along the z-axis, we introduce XXZ spin anisotropy into either the intrachain or interchain couplings, a feature known to exist, albeit moderately, in real materials. The magnetisation per site, serving as an order parameter for the CAFO, is defined by:
$m=\frac{1}{l_{x}l_{y}}\left\lvert\sum_{\alpha,j}(-1)^{j}\langle S^{z}_{\alpha,j}\rangle\right\lvert$ (6)
Given the bipartite lattice configuration perpendicular to the chains, we segregate the chain indices $j$ into even and odd numbers for each sublattice. For a comprehensive exploration of parameter dependencies, we confine our system to 2-chain cluster ($l_{y}=2$), a scenario that maximises quantum fluctuations, thereby posing a challenging environment for the establishment of the CAFO. Nonetheless, the formation of a CAFO under such stringent conditions suggests that more stable CAFO structures would naturally arise in systems with reduced quantum fluctuations, such as in two- or three-dimensional lattices. Indeed, we have confirmed a more pronounced CAFO stabilisation for 4-chain ($l_{y}=4$) and 8-chain ($l_{y}=8$) clusters illustrated in Fig. 1f in comparison to the $l_{y}=2$ case.
Using the DMRG method, we calculate the magnetisation per site in the thermodynamic limit ($l_{x}\to\infty$) for various strengths of the XXZ anisotropy and representative values of the ratio $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$. The results are plotted as a function of the interchain couplings $J_{\rm a}$ and $J_{\rm b}$ in Fig. 3. In Fig. 3a, the XXZ anisotropy ($\Delta_{1}<1$) is introduced in the intrachain coupling, while in Fig. 3b, the XXZ anisotropy ($\Delta_{\rm a,b}<1$) is included in the interchain coupling. In both cases, a critical interchain coupling exists for all combinations of $J_{2}/\lvert J_{1}\rvert$ and a kind of the interchain couplings, beyond which the system transitions to the CAFO phase. When $J_{1}$-$J_{2}$ chains in the spiral phase are coupled by frustrated AFM interchain couplings, increasing the interchain coupling reduces the propagation number. Once this number reaches zero, the system transitions to the CAFO phase [43]. Therefore, comparing the results for $J_{2}/\lvert J_{1}\rvert=1/3$ and $J_{2}/\lvert J_{1}\rvert=1/2$, the interchain coupling required to induce the transition to the CAFO phase is smaller for $J_{2}/\lvert J_{1}\rvert=1/3$, which originally has a smaller propagation number. Furthermore, as inferred from Eq. 4, the effect of $J_{\rm b}$ is generally larger than that of $J_{\rm a}$. For instance, for $J_{2}/\lvert J_{1}\rvert=1/3$, a slight XXZ anisotropy in the intrachain coupling leads to a significant magnetisation when the interchain coupling is $J_{\rm b}$. This scenario corresponds to the case of Li2CuO2 discussed in Sec. 4.1.
Thus, it was found that each $J_{1}$-$J_{2}$ chain becomes magnetised by the AFM interchain coupling as if an external uniform field had been applied. In the following subsection, we will verify whether MBS indeed emerges in this magnetised state.
3.3 Magnon excitation spectrum
Figure 4: (a) Schematic illustration distinguishing a MBS from a non-MBS within the MES. (b,c) DDMRG results for the MES $G_{p}(\omega)$ at various values of $p=1$, $2$, $3$, and $4$, for representative ratios $J_{2}/\lvert J_{1}\rvert=1/3$ and $1/2$, with the XXZ anisotropy applied to (b) inchain and (c) interchain couplings. (d) Angle-resolved MES $G_{1}(k,\omega)$ and $G_{2}(k,\omega)$ for $J_{2}/\lvert J_{1}\rvert=1/2$, $J_{\rm a}/\lvert J_{1}\rvert=1.0$, and $\Delta_{\rm a}=0.1$.
In real materials, the behaviour of coupled $J_{1}$-$J_{2}$ chains can be particularly fascinating when considering MBSs. A straightforward approach to detect these states is to calculate the binding energy of magnons, $E_{\rm b}(M,p)$, as done in Sec. 2.2 using Eq. (1). However, this approach faces a problem for coupled chains due to interchain exchange interactions, which lead to the non-conservation of the total $S^{z}$ within each chain. Consequently, Eq. (1) cannot be directly applied. To circumvent this issue, we investigate the presence of MBSs by calculating the magnon excitation spectrum (MES) for scenarios where $p$ spins are flipped simultaneously. The MES for $p$-magnon flips is defined as:
$G_{p}(\omega)=\sum_{n>0}\lvert\langle n\lvert\prod_{i=r}^{r+p-1}S^{-}_{i}\lvert 0\rangle\lvert^{2}\delta(\omega-E_{n}+E_{0}).$ (7)
This function is crucial for identifying MBSs. It has been established that when $p$ magnons form a bound state, the average size of these magnons is effectively $p-1$ [13]. This implies that the magnons are tightly bound and form the lowest-lying excitations in the energy spectra. Therefore, computing MES by flipping $p$ contiguous spins is a reasonable approach to explore the dynamics and interactions of MBSs in coupled chains. This methodology provides a deeper understanding of quantum magnetic phenomena and extends our insight into the collective excitations in complex spin systems.
To determine the presence of MBSs, one must first compute the MES for $p=1$, denoted as $G_{1}(\omega)$, and for $p>1$, denoted as $G_{p(>1)}(\omega)$. As illustrated in the left panel of Fig. 4a, if the onset of $G_{1}(\omega)$ occurs at a lower energy than that of $G_{p(>1)}(\omega)$, this suggests that there is no binding among the magnons. Conversely, as shown in the right panel of Fig. 4a, if the onset of $G_{p(>1)}(\omega)$ is at a lower energy than that of $G_{1}(\omega)$, it indicates that $p$ magnons are bound. In this scenario, the difference between the onsets of $G_{p(>1)}(\omega)$ and $G_{1}(\omega)$ corresponds to the magnon binding energy $E_{\rm b}(M,p)$. To accurately identify the state with the lowest energy onset, one needs to systematically vary $p$ through values such as $2$, $3$, $4$, and so on, and observe which $G_{p(>1)}(\omega)$ yields the minimum onset energy. This procedure enables the identification of the optimal number of bound magnons and provides a quantitative measure of the magnon binding energy. It is noteworthy that when MBSs are not formed, the system is essentially gapless. Consequently, the lowest excitation of $G_{1}(\omega)$, corresponding to inelastic neutron scattering (INS), is at zero energy.
We then investigate whether MBSs truly form when a CAFO is established by the internal field, as demonstrated in Sec. 3.2. In the CAFO, each $J_{1}$-$J_{2}$ chain orders ferromagnetically, with neighbouring $J_{1}$-$J_{2}$ chains having their magnetisations aligned anti-parallel along the $z$-axis. Here, we compute the MES for the $J_{1}$-$J_{2}$ chains with magnetisation pointing in the positive $z$-direction using Eq. (7). We note that an equivalent result is obtained for chains with magnetisation in the negative $z$-direction by replacing $S^{-}_{i}$ with $S^{+}_{i}$ in Eq. (7).
The MES computed using DDMRG method for various parameters is shown in Figs. 4b and 4c. The parameters used in those calculations are marked with circles in Fig. 3. To clarify the impact of CAFO on the MES, we compare the MES before and after the critical interchain coupling. In Fig. 4b, where $J_{1}$ is modified to include XXZ anisotropy, we observe that for $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{1}=0.9$, no magnon binding is present at $J_{\rm b}=0.05$, as the lowest excitation is provided by $G_{1}(\omega)$. However, when $J_{\rm b}=0.1$, leading to the formation of CAFO, $G_{2}(\omega)$ provides the lowest excitation, indicating that the ground state is a nematic state. Intriguingly, the onset of $G_{3}(\omega)$ is very close to that of $G_{2}(\omega)$, suggesting that a triatic state exists very near the ground state. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{1}=0.8$, the MES transitions from $G_{1}(\omega)$ to $G_{2}(\omega)$ as the lowest excitation when CAFO forms.
Fig. 4c shows the MES when XXZ anisotropy is introduced in the interchain coupling. Essentially, similar results are obtained as when XXZ anisotropy is applied to $J_{1}$. For $J_{2}/\lvert J_{1}\rvert=1/3$ and $\Delta_{\rm a}=0.1$, the ground state becomes a triatic state upon the formation of CAFO. For $J_{2}/\lvert J_{1}\rvert=1/2$ and $\Delta_{\rm a}=0.1$, the ground state becomes a nematic state upon CAFO. The value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$. Given that $J_{2}/\lvert J_{1}\rvert=1/3$ is very close to the boundary between $p=2$ and $p=3$ at $J_{2}/\lvert J_{1}\rvert=0.367615$ in the high-field phase diagram of a single $J_{1}$-$J_{2}$ chain (see Figs. 2b, c, d), the MES shows $G_{2}(\omega)$ and $G_{3}(\omega)$ onsets close to each other, almost degenerate at the ground state. However, in any case, the presence of CAFO results in excitations significantly lower than those given by $G_{1}(\omega)$.
In comparison with the properties of conventional MBSs, it would also be useful to examine the total momentum of the multi-magnon modes within the MBS formed by the internal field. To confirm this, we compute the angle-resolved MES, which can be defined as follows:
$G_{p}(k,\omega)=\sum_{n>0}\lvert\langle n\lvert\sum_{j}\left(\prod_{i=j}^{j+p-1}S^{-}_{i}\right)\exp(ikr_{j})\lvert 0\rangle\lvert^{2}\delta(\omega-E_{n}+E_{0})$ (8)
The results of calculating $G_{p}(k,\omega)$ using DDMRG for $G_{1}(\omega)$ and $G_{2}(\omega)$ corresponding to $J_{2}/\lvert J_{1}\rvert=1/2$, $J_{\rm a}/\lvert J_{1}\rvert=1.0$, $\Delta_{\rm a}=0.1$ are presented in Fig. 4d. Essentially, these results demonstrate that our novel MBS exhibits properties akin to those emerging in an isolated $J_{1}$-$J_{2}$ chain under an external magnetic field [10, 11, 12, 13, 14, 15]. Specifically, the two-magnon excitation $G_{2}(k,\omega)$ shows a minimum at momentum $k=\pi$. Conversely, $G_{1}(k,\omega)$, which corresponds primarily to the transverse component of the dynamical spin structure factor $S^{xx}(k,\omega)$ (or $S^{yy}(k,\omega)$), exhibits a minimum around $k=0.332\pi$, reflecting the incommensurate correlations of the isolated $J_{1}$-$J_{2}$ chain in the absence of a magnetic field. Given the nematic ground state, the lowest bound of $G_{2}(k,\omega)$ is located lower than that of $G_{1}(k,\omega)$, consistent with the $G_{1}(\omega)$ and $G_{2}(\omega)$ shown in Fig. 4c.
4 Realisation of zero-field MBS in Q1D cuprates
Finally, we consider four materials as candidates for the potential formation of MBSs based on the mechanism proposed above.
Figure 5: (a) DDMRG results for $G_{p}(\omega)$ ($p=1$ to $4$) using Li2CuO2 parameter set with $12\times 8$-chain cluster. (b) Intensity map of the INS data extracted from Ref. [44]. (c,d) DDMRG results for $G_{p}(\omega)$ ($p=1$ to $4$) using the intrachain couplings for LiCuSbO4 and PbCuSO4(OH)2 with $24\times 4$-chain cluster. The interchain AFM couplings are set to be tunable parameter.
4.1 Li2CuO2
The most promising material for the application of our MBS emergence mechanism is Li2CuO2. The magnetic couplings for this material are estimated to be $J_{1}=-228$ K, $J_{2}=75.7$ K, $J_{\rm a}=1$ K, and $J_{\rm b}=9$[44]. Also, paramagnetic and AFM resonance experiments have observed easy-axis anisotropy along the a-axis [45], which is estimated to be approximately $5\%$ of $J_{1}$ [46, 44, 47]. Therefore, we assume a small $J_{1}$ XXZ anisotropy with $\Delta_{1}=0.95$. Given that $J_{2}/\lvert J_{1}\rvert=0.332(>1/4)$, each $J_{1}$-$J_{2}$ chain would remain in the spiral phase if the interchain couplings were absent. However, the introduction of small interchain couplings $J_{\rm a}=1$ K and $J_{\rm b}=9$ K results in the system exhibiting a CAFO. It has been estimated that a $J_{\rm b}$ of $8.2$ K is required to induce CAFO [48], and the actual interchain coupling slightly exceeds this value. As shown in Fig. 1b, this system has a three-dimensional structure where each chain is coupled to four neighbouring chains through $J_{\rm a}$ and $J_{\rm b}$ (see Fig. 1d). Therefore, we perform our DMRG calculations using a $12\times 8$-chain cluster, applying open boundary conditions along the chain direction and periodic boundary conditions perpendicular to the chains, as depicted in Fig. 1f. Using the aforementioned parameters, we confirm a CAFO state with $2\lvert(\rvert\langle S^{z}_{i}\rangle)=0.965$. This magnitude closely matches the experimentally estimated Cu ion moment of $0.96(4)\mu_{\rm B}$ at $T=1.5$[49], further supported by ab-initio calculations [50, 51, 46].
We then examine the results of the MES calculations using DDMRG method. Fig. 5a plots $G_{p}(\omega)$ for $p$ values from 1 to 4. Focusing on a single $J_{1}$-$J_{2}$ chain, the FM order with a large magnetization of $0.9652\lvert\langle S^{z}_{i}\rangle\rvert=$ 0.965 resembles a state almost fully saturated by an external magnetic field. Consequently, one might expect an MBS to form. The spectra clearly show that the $p=3$ spectrum $G_{3}(\omega)$ has the lowest excitation at $\omega\sim 2.58$ meV, indicating a triatic ground state. The energy gap between this peak and the lowest energy peak of $G_{1}(\omega)$ at $\omega\sim 2.73$ meV is $\Delta_{\rm gap,1}\sim 0.15$ meV, corresponding to the magnon binding energy. Note that our DDMRG calculations are performed at $T=0$. Additionally, $G_{1}(\omega)$ exhibits another gap, $\Delta_{\rm gap,2}\sim 1.41$ meV, between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV, likely attributed to the spin gap induced by XXZ anisotropy. However, since the ground state is not a simple singlet, $\Delta_{\rm gap,2}$ is not a perfect gap, and some in-gap states are present between $\omega\sim 2.73$ meV and $\omega\sim 4.17$ meV.
For comparison, Fig. 5b shows the experimental INS results near $k=0$. Given that the ground state ($\omega=0$) is a triatic state, the lowest energy peak of $G_{3}(\omega)$ corresponds to $\Delta E=0$. Meanwhile, $G_{1}(\omega)$ corresponds to the structure factor $S^{xx}(k,\omega)$ (or $S^{yy}(k,\omega)$), which is observable in INS experiments. Although the resolution around $\Delta E=0$ is limited in this INS data, a nearly-flat dispersion appears near $\Delta E=0.12$ meV (also see Fig. 2(c) in Ref. [44]), corresponding to our calculated $\Delta_{\rm gap,1}$ of $0.15$ meV. Furthermore, a gap of approximately $1.28$ meV exists from this nearly-flat dispersion to the lower bound of the FM dispersion, aligning with our calculated $\Delta_{\rm gap,2}$ and suggesting the presence of in-gap state weights. Note that this INS observation was made at $T=4.1$ K, while the critical temperature of the triatic state is estimated to be around $T_{\rm c}=1$ K by assuming the BCS prediction between energy gap at $T=0$ and the critical temperature, $\Delta(T=0)=1.76k_{\rm B}T_{\rm c}$ [52]. Possibly, the gapped excitation, i.e., the lowest excitation peak $G_{1}(\omega)$ near $\Delta E=0.12$ meV, might be rather broadened. Therefore, more definitive evidence of MBS formation would require lower temperature and higher resolution experiments. Additionally, as demonstrated with the materials listed below, applying pressure to strengthen the CAFO could increase the MBS binding energy, potentially enhancing $\Delta_{\rm gap,1}$, observable through experimental measurements.
4.2 Ca2Y2Cu5O10
We next consider the compound Ca2Y2Cu5O10 [53]. Its crystal structure is illustrated in Fig. 1c. This material exhibits a CAFO below $T_{\rm N}=29.5$ K, where spins align ferromagnetically along the a-axis (chain direction) and antiferromagnetically along the c-axis [54, 55]. The ordered magnetic moment is $0.9\mu_{\rm B}$. While the spins also align ferromagnetically along the b-axis, the interlayer couplings in the b direction are negligible compared to those within the ac layer. Thus, the ferromagnetic stacking of 2D CAFO planes can be regarded as the ”2D analogue” of Li2CuO2.
However, a key difference from Li2CuO2 in the context of our MBS discussion is that the in-chain couplings in the $J_{1}$-$J_{2}$ chain for Ca2Y2Cu5O10 are $J_{1}=-24$ meV and $J_{2}=5.5$ meV, leading to $J_{2}/\lvert J_{1}\rvert=0.23(<0.25)$ [56]. This implies that even in the absence of interchain couplings, each $J_{1}$-$J_{2}$ chain would order ferromagnetically. The CAFO in this material is driven by AFM interchain couplings, estimated as $J_{\rm a}+J_{\rm a}\sim 2.29$ meV [57, 56]. Although the FM chains have their magnetisation directions fixed by the internal field, MBS states do not emerge.
Nevertheless, a notable point is that $J_{2}/\lvert J_{1}\rvert=0.23$ is very close to the critical point of $0.25$. Therefore, tuning $J_{2}/\lvert J_{1}\rvert$ to exceed $0.25$ may not be unrealistic. For instance, applying pressure might achieve this. Within the effective spin Hamiltonian, the smaller couplings are expected to be most sensitive to pressure. Consequently, it can be anticipated that the change in $J_{2}$ would be more significant than that in $J_{1}$ under pressure, making it feasible to achieve $J_{2}/\lvert J_{1}\rvert>0.25$. Once $J_{2}/\lvert J_{1}\rvert>0.25$ is realized, the CAFO could be retained by the relatively large interchain couplings $J_{\rm a}+J_{\rm a}\sim 2.29$ meV, potentially leading to the observation of MBS with a relatively large binding energy. Furthermore, if the value of $J_{2}/\lvert J_{1}\rvert$ can be continuously varied by pressure, it might be possible to observe MBS ground state that successively changes $p$ with pressure.
Nevertheless, a notable point is that $J_{2}/\lvert J_{1}\rvert=0.23$ is very close to the critical point of $0.25$. Therefore, tuning $J_{2}/\lvert J_{1}\rvert$ to exceed $0.25$ may not be unrealistic. For instance, applying pressure/stretch might achieve this. Within the effective spin Hamiltonian, the smaller couplings are expected to be most sensitive to pressure. Consequently, it can be anticipated that the change in $J_{2}$ would be more significant than that in $J_{1}$ under pressure/stretch, making it feasible to achieve $J_{2}/\lvert J_{1}\rvert>0.25$. Once $J_{2}/\lvert J_{1}\rvert>0.25$ is realised, the CAFO could be retained by the relatively large interchain couplings $J_{\rm a}+J_{\rm a}\sim 2.29$ meV, potentially leading to the observation of MBS with a relatively large binding energy. Furthermore, if the value of $J_{2}/\lvert J_{1}\rvert$ can be continuously varied by pressure, it might be possible to observe an MBS ground state that successively changes $p$ with pressure. Actually, a possible tuning of $J_{2}/\lvert J_{1}\rvert$ by hydrostatic pressure has been suggested in Ref. [58].
4.3 LiCuSbO4
The third material under consideration is LiCuSbO4 [59]. In this compound, the magnetic couplings between layers are negligible, allowing it to be effectively treated as a 2D system. The structure of the interchain couplings is equivalent to that depicted in Fig. 1e. The intrachain couplings have been estimated as $J_{1}=-125$ K and $J_{2}=37.6$[23], leading to a ratio of $J_{2}/\lvert J_{1}\rvert=0.3$. Given this ratio is very close to the critical value of 0.25, even moderately strong AFM interchain couplings could stabilise a CAFO state. However, in practice, LiCuSbO4 exhibits short-range incommensurate spin correlations below $T\sim 9$ K and, unlike the spiral spin-chain compounds LiCuVO4 [60, 61] and LiCuZrO4 [62, 63], does not show long-range magnetic order down to $T\sim 0.1$ K. This may suggest that the interchain couplings in LiCuSbO4 are very weak. Indeed, DFT and DFT+$U$ band structure calculations estimate $J_{\rm d}\sim 1$[23].
If the interchain couplings can be slightly enhanced through the application of pressure, CAFO might be realised, potentially leading to the manifestation of MBS. To explore this, we vary $J_{\rm d}$ around 1 K and calculate the MES $G_{p}(\omega)$ for a $24\times 4$-chain cluster. Assuming an exchange anisotropy of approximately $10\%$ [59, 23, 64], we set $\Delta_{1}=0.9$. The DDMRG results for $G_{p}(\omega)$ with $J_{\rm d}=0.75$ K and $1.25$ K, for $p$ values from 1 to 4, are shown in Fig. 5c.
For $J_{\rm d}=0.75$ K, CAFO is not yet stabilised, and $G_{1}(\omega)$ exhibits the lowest energy excitation, indicating the absence of magnon binding. Nonetheless, the proximity of the lowest energy excitation in $G_{3}(\omega)$ suggests that a slight increase in $J_{\rm d}$ might stabilise a triatic ground state. Then, examining the results for $J_{\rm d}=1.25$ K, it is evident that $G_{3}(\omega)$ exhibits the lowest energy excitation, with a relatively large magnon binding energy of approximately 1 meV. Given that LiCuSbO4 is already near the CAFO phase with $J_{2}/\lvert J_{1}\rvert=0.3$, only a modest application of pressure might be required to observe both CAFO and MBS. Interestingly, as $J_{\rm d}$ increases, $G_{1}(\omega)$ shifts to higher energies while changing shape, whereas the spectra for $p>1$ remain largely unchanged. This observation supports the statement that the value of $p$ for the ground state generally corresponds to the number of magnon bindings stabilised by an external field in a single $J_{1}$-$J_{2}$ chain for a given $J_{2}/\lvert J_{1}\rvert$.
4.4 PbCuSO4(OH)2
The forth material we discuss is PbCuSO4(OH)2, commonly known as linarite [65], which has a 2D layer lattice structure with interchain coupling $J_{\rm d}$, as illustrated in Fig. 1e. The in-chain couplings for this material have been estimated as $J_{1}=-100$ K and $J_{2}=36$[66], with a ratio of $J_{2}/\lvert J_{1}\rvert=0.36$, which is greater than the FM critical point of 0.25. Unlike LiCuSbO4, linarite exhibits long-range spiral order with a small propagation number $q=0.186\pi$ along the $J_{1}$-$J_{2}$ chain below $T_{\rm N}=2.8$ K. This long-range order is believed to be stabilized by a sizable AFM interchain coupling, estimated to be $J_{\rm d}\sim 6$[67]. However, for this material to exhibit CAFO, an interchain coupling of approximately $J_{\rm d}=0.1\lvert J_{1}\rvert\sim 10$ K is required. Similar to LiCuSbO4, there is potential for pressure effects to enhance the interchain coupling. If $J_{\rm d}>10$ K can be achieved, we can explore the possibility of realizing an MBS ground state.
Given the highly anisotropic nature of PbCuSO4(OH)2, we set $\Delta_{1}=0.9$, though the actual anisotropy may be slightly larger. Fig. 5c plots the MES $G_{p}(\omega)$ obtained from DDMRG calculations on a $24\times 4$-chain cluster with $J_{\rm d}=15$ K. The lowest energy excitation state near $\omega=1.3$ meV is given by $G_{2}(\omega)$, indicating that the ground state is a nematic state. This observation aligns with the realization of a nematic state in an external magnetic field for $J_{2}/\lvert J_{1}\rvert=0.36$.
4.5 Other materials
Other materials that potentially apply to our MBS emergence mechanism include CuSb2O4 [68] and Rb2Cu2Mo3O12 [69], which as also modeled by coupled FM-AFM $J_{1}$-$J_{2}$ chains. The ratios of their intrachain couplings are estimated to be $J_{2}/\lvert J_{1}\rvert\sim 0.5$ and $J_{2}/\lvert J_{1}\rvert=0.37$, respectively. Due to the relatively small interchain couplings, these materials are thought to exhibit a spiral ordered state at low temperatures. As with the previously discussed materials, if these can transition to a CAFO state under the influence of pressure, MBS might be realised.
Another material worth mentioning is LiVCuO4, frequently discussed as a candidate material for MBS under high external fields [16, 18]. This material has extremely small interchain couplings, making it ideal for observing MBS at high magnetic fields [70]. The ratio of intrachain couplings has not reached a complete consensus, with estimates ranging from $J_{2}/\lvert J_{1}\rvert=0.75$ [70, 43], to 1.42 [71], and 3.49 [61]. In any case, these values are significantly distant from the FM critical point $J_{2}/\lvert J_{1}\rvert=0.25$, indicating that substantial interchain couplings would be necessary to induce a transition to CAFO. Therefore, a pressure-induced transition to CAFO might be unrealistic, and thus LiVCuO4 can be eliminated from our list of candidate materials for our MBS mechanism.
Until now, we have considered systems where $J_{1}$-$J_{2}$ chains with a bipartite structure transition to CAFO under AFM interchain coupling. However, a simple bipartite structure is not that necessary. For instance, the Q1D mineral antlerite Cu3SO4(OH)4. The magnetic model of this system can be represented as three-leg zigzag ladders, where the central leg hosts competing AFM exchanges, and each side leg features alternating NN FM exchanges $J_{1}$, $J_{1}^{\prime}$ and a weaker NNN AFM exchange $J_{2}$. These three chains are coupled by zigzag-shaped AFM interchain couplings. The side chains are dimerised FM-AFM $J_{1}$-$J_{2}$ chains, with parameters estimated as $J_{1}=-26$, $J_{1}^{\prime}=-11$, and $J_{2}=-6$ (in meV) by GGA+U approximation. Without interchain coupling, the side chains would be in a spiral state [72, 73]. However, the experimentally observed low-temperature magnetic structure shows that the side legs are ferromagnetically polarised in opposite directions, while the central chain is in an AFM order. This suggests that the effect of interchain coupling causes the side chains to polarise ferromagnetically.
There is a possibility that MBS emerges due to the effect of interchain couplings in the side chains. Materials with such structures are diverse, indicating a wide potential for observing zero-field MBS, which has not been considered previously. We anticipate that our proposed mechanism for the emergence of MBSs will be of significant interest to experimentalists.
5 Conclusion
In this study, we have demonstrated a novel mechanism for the stabilization of MBS in Q1D cuprates through interchain interactions. Unlike conventional MBS that require an external magnetic field, our results show that AFM interchain coupling can induce MBS even in the absence of an external field. This finding is significant as it provides a new perspective on the interplay between interchain and intrachain interactions in low-dimensional magnetic systems.
Our numerical simulations using the DMRG method confirm that the AFM interchain coupling can act similarly to an internal magnetic field, leading to the formation of CAFO and the subsequent stabilization of MBS. This is particularly intriguing because the AFM interchain coupling has a dual role; while it destabilizes MBS at high fields, it stabilizes them at zero field.
Notably, we found that MBS can be stabilized in Li2CuO2 under ambient pressure and zero magnetic field, highlighting the robustness of this mechanism. However, this aspect has not been experimentally investigated, as it has not been viewed from this perspective before. Additionally, we explored the conditions under which this phenomenon occurs in other materials such as Ca2Y2Cu5O10, LiCuSbO4, and PbCuSO4(OH)2. Our findings suggest that the application of pressure or chemical substitution could further enhance the interchain coupling in these materials, making it feasible to experimentally observe the MBS predicted by our theoretical models.
Revisiting various physical phenomena from the perspective of effective internal magnetic fields due to interchain coupling may lead to new discoveries. This study serves as a promising example of such potential, demonstrating the significant implications for understanding magnetic excitations in Q1D systems.
Methods
Density-matrix renormalization group
We employ the density-matrix renormalization group (DMRG) technique [74] to study the ground-state properties. For single-chain calculations, we study periodic chains with lengths up to $l_{x}=192$ keeping $\chi=800$ density-matrix eigenstates for renormalization procedure. The discarded weight is negligible. When the magnon binding energy is calculated, it is better to use periodic chains to avoid picking up low excitations that may occur near the edges of open chains. For two-chain calculations, we study open ladders with lengths up to $l_{x}=280$ keeping $\chi=3600$. The largest discarded weight is $w_{\rm d}\sim 10^{-11}$. To explicitly fix the direction of magnetization, which corresponds to the direction in which spin-rotation symmetry is broken in each chain, magnetic fields are applied to the sites at both ends of an open ladder. The direction of the magnetic fields is parallel to the z-axis, with fields at both ends of a single chain aligned in the same direction, and those on adjacent chains in opposite directions. However, the magnitude of magnetization extrapolated to the thermodynamic limit is independent of the values of these edge fields. For $12\times 8$-chain calculations, we keep $\chi=1200$ leading to $w_{\rm d}\sim 10^{-10}$. In this case, we focus on the collinear antiferromagnetic ordered state, where the quantum fluctuations are significantly suppressed, the DMRG accuracy is very good. Similar to the case of the open ladder, magnetic fields are applied to both ends of the system.
Dynamical density-matrix renormalization group
We employ the dynamical density-matrix renormalization group (DDMRG) technique [75] to calculate the magnon excitation spectrum (MES), defined as We calculate the spin excitation spectrum
$G_{p}(k,\omega)=\frac{1}{\pi}{\rm Im}\langle 0|B^{+}_{-k}\frac{1}{\hat{H}+\omega-E_{0}-{\rm i}\eta}B^{-}_{k}|0\rangle\,,$ (9)
where $B^{-}_{k}=\sum_{j}\prod_{i=j}^{j+p-1}S^{-}_{i}\exp(ikr_{j})$ for $p$-magnon flips, $|0\rangle$ and $E_{0}$ are, respectively, the wavefunction and energy of the ground state of the Hamiltonian (3). A Lorentz distribution of width $\eta$ is used to broaden the delta peaks at the excited states. Since the DDMRG algorithm performs best for open boundary conditions, we study open clusters along the chain direction. When we calculate the local MES $G_{p}(\omega)=\sum_{k}G_{p}(k,\omega)$, the site position is fixed at $j=l_{x}/2$.
In two-chain calculations, we examine ladders of length $l_{x}=32$ with the broadening $\delta(\omega/p)=0.04\lvert J_{1}\rvert$. The DDMRG approach is based on a variational principle, therefore it is necessary to prepare a ’good trial function’ of the ground state with the density matrix eigenstates. We keep $\chi_{1}=1200$ in the first few dozen DMRG sweeps to obtain the ground state, and then to keep $\chi_{2}=400$ to calculate the excitation spectrum. This approach ensures that the maximum discarded weight is of the order of $10^{-6}$. For $24\times 4$-chain calculations, the following parameters were set: $\chi_{1}=2000$, $\chi_{2}=600$, and $\eta/p=0.0025\lvert J_{1}\rvert$, which resulted in a discarded weight of approximately $10^{-5}$. For $12\times 8$-chain calculations, the following values were set: $\chi_{1}=1600$, $\chi_{2}=600$, and $\delta(\omega/p)=0.0025\lvert J_{1}\rvert$. This resulted in a value of $w_{\rm d}\sim 10^{-7}$.
[…]
Supplementary information
Finite-size effects in the MES are discussed in Supplemental Information.
[…]
Acknowledgments
We thank Ulrike Nitzsche for technical assistance. We thank S.D.Wilson and R.O.Kuzian for fruitful discussion. This project is funded by the German Research Foundation (DFG) via the projects A05 of the Collaborative Research Center SFB 1143 (project-id 247310070).
Authors’ contributions C.E.A. and S.N. performed numerical calculations and analyzed the results. S.N. designed the study and drafted the manuscript in discussion with C.E.A. and S.-L.D.. All authors approved the final manuscript.
Competing financial interests The authors declare no competing interests.